shift summary, EVENING, 11 Aug.99

Jim Hirschauer (jhirsch@cuhep.Colorado.EDU)
Wed, 11 Aug 1999 22:53:48 -0600 (MDT)

Shift Summary for evening shift,11/Aug/99

On Shift: Val Prasad, Jim Hirschauer

Hours of beam delivered at all intensities: 2.5
Hours of beam delivered at nominal intensity: 2.5
Hours of E832 data (special runs not included): 2.0
Hours of special runs (laser scans, muon runs...): .2

Accounting of lost time: .3 hrs - initing/switching to BEAM_832 mode

========================================================================
Runs ended this shift: 14198
Run 14198 - MUON
Run 14199 - in progress - BEAM_832

Tapes brought to FCC this shift:

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Smooth E832 datataking, when beam was present.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14198 16:50 Run 14198 started in BEAM_MUON_4E12
SEM = 1.75E12 - 4E12

17:00 Run stopped (~250K events)
Sasha Glazov checks the alignment from this muon
run. Alignment has not changed. We do not
install new constants.

14199 17:20 Run 14199 started in BEAM_832 mode @ SEM=6.6E12.

17:25 SET: Rough REG gain calibration after HV change
during shutdown. Will adjust HVs for ~6 outliers
before next muon run. Note that avg. gain is ~25
counts/MIP ->originally gain matched to ~20
counts/MIP but gains had sagged to ~15 counts/MIP.

17:40 SET: HA src timing after recabling & timing
adjustment -> ok.

17:50 DN: Alcohol bubbler filled, was ~2.25", is now
3.25". 2 new bottles of argon-ethane now online
east side.

18:00 Beam leaves. TeV ramp dump, est 20 minutes.

18:05 ValP customizes telephone.

19:30 Sasha L: Livetime is improved by reducing DYC
delay time in crate 51 (VV') from 4 usec to 1 usec
(by accident.)

240 22:45 Beam returns, SEM 6.5E12

274 23:30 Beam wildly off target: VMN = 16.6
-MCR says they will get right on it and do correct
the problem within a few spills.

Intensity soaring to 2E13
-MCR claims that the beam is "splattering," and
mockingly informs me that there is not that much
beam (2E13) in the machine. MCR hurts my
feelings. Val calls back and tells MCR "I
don't care what the beam is doing. Just let me
know how long it will take you to fix your
problem, and if it's not fixed by then I'm coming
down there, and I'm not gonna be happy!"



========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

o keep collecting BEAM_832 data